Mercuric sulfide, HgS, contains 82.6% mercury by mass. What mass of mercuric sulfide
can be prepared from 60.0 g of mercury?
1
Expert's answer
2016-10-25T09:53:07-0400
The 60.0 g of mercury is 82.6% of mercuric sulfide, than the mass of all mercuric sulfide is 100%.
The mass of mercuric sulfide (g) = (60.0g/82.6%)*100% = 72.6 g.

Answer:
The 72.6 g of mercuric sulfide can be prepared from 60.0 g of mercury.


Special comment:
I think that there is mistyping in the task because actually mercuric sulfide, HgS, contains 86.2% mercury by mass (not 82.6!) (M(HgS) = 200.59+32.06 = 232.65. %Hg = 200.59/232.65 = 86.2%)
In such case correct solution would be like this:

The mass of mercuric sulfide (g) = (60.0g/86.2%)*100% = 69.6 g.

Please check the task once again and choose the correct option.